OpenStudy (anonymous):

Which of the following equations has only one solution? 1) x2+ 4x + 4 = 0 2) x2+ x = 0 3) x2- 1 = 0

OpenStudy (blurbendy):

What do you get when you factor the first one?

OpenStudy (blurbendy):

@Gabriella77

OpenStudy (anonymous):

x2+4x+4=0 >> 6x = -4 >> x =-4/6

OpenStudy (blurbendy):

no, the expression is x^2 + 4x + 4 = 0 not 2x + 4x + 4 = 0 see the difference?